# Efficient Adjusted Joint Significance Test and Sobel-Type

### v7

@@ 5.3 Coverage Probability of Confidence Interval
The performance of the ASobel-type confidence interval presented in Section 4 is evaluated through simulations conducted in this subsection. The traditional Sobel-type confidence interval, ${\rm CI_{Sobel}}$, given in (4.2), is also considered for comparison. The data are generated in a similar manner as those described in Section 5.2, with the parameters chosen as (i) Linear mediation model : ${\bm{\alpha}}=(0,0.35,0.5,0,0,0.15,0.25)^{\prime}$ and ${\bm{\beta}}=(0,0,0,0.35,0.5,0.3,0.35)^{\prime}$; (ii) Logistic mediation model: ${\bm{\alpha}}=(0,0.35,0.5,0,0,0.25,0.35)^{\prime}$ and ${\bm{\beta}}=(0,0,0,0.55,0.75,0.25,0.35)^{\prime}$; (iii) Cox mediation model: ${\bm{\alpha}}=(0,0.35,0.5,0,0,0.25,0.35)^{\prime}$ and ${\bm{\beta}}=(0,0,0,0.5,0.45,0.25,0.35)^{\prime}$. The sample sizes are chosen as $n=500$ and 1500, respectively. All the results are based on 5000 repetitions.
The coverage probability (CP) and length of the 95% confidence interval (LCI) with ${\rm CI_{Sobel}}$ and ${\rm CI_{ASobel}}$ are reported in Tables 3, S.5 and S.6. The ${\rm CI_{ASobel}}$ outperforms ${\rm CI_{Sobel}}$ in terms of CP under $\alpha=\beta=0$. Additionally, the LCI of ${\rm CI_{ASobel}}$ is significantly shorter compared to that of ${\rm CI_{Sobel}}$ when $\alpha\beta=0$. These findings are consistent with the result of Theorem 3.
@@
Table 4: The comparison between AJS and AB-JS in terms of size and ${}^{\ddagger}$.
@@ 6 Real Data Examples
Dataset I: (continuous outcomes). The Louisiana State University Health Sciences Center has explored the relationship between children weight and behavior through a survey of children, teachers and parents in Grenada. The dataset is publicly available within the R package mma. To perform mediation analysis as that of Yu and Li (2017), we set gender as the exposure $X$ (Male =0; Female = 1), and the outcome $Y$ is body mass index (BMI). We consider three mediators: $M_{1}$ (join in a sport team or not), $M_{2}$ (number of hours of exercises per week) and $M_{3}$ (number of hours of sweating activities per week). Furthermore, there are three covariates $Z_{1}$ (age), $Z_{2}$ (number of people in family) and $Z_{3}$ (the number of cars in family). After removing those individuals with missing data, we totally have 646 samples when conducting mediation analysis in the context of logistic mediation models. We consider the logistic mediation model to fit this dataset:
@@
Dataset III: (time-to-event outcomes). We apply our proposed method to a dataset from The Cancer Genome Atlas (TCGA) lung cancer cohort study, where the data are freely available at https://xenabrowser.net/datapages/. There are 593 patients with non-missing clinical and epigenetic information. From Luo et al. (2020) and Zhang et al. (2021b), we use seven DNA methylation markers as potential mediators: $M_{1}$ (cg02178957), $M_{2}$ (cg08108679), $M_{3}$ (cg21926276), $M_{4}$ (cg26387355), $M_{5}$ (cg24200525), $M_{6}$ (cg07690349) and $M_{7}$ (cg26478297). The exposure $X$ is defined as the number of packs smoked per years, and the survival time is the outcome variable. Two hundred forty three patients died during the follow-up, and the censoring rate is 59%. We are interested in testing the mediation effects of DNA methylation markers along the pathways from smoking to survival of lung cancer patients. Four covariates are included: $Z_{1}$ (age at initial diagnosis), $Z_{2}$ (gender; male = 1, female=0), $Z_{3}$ (tumor stage; Stage I = 1, Stage II = 2, Stage III = 3, Stage IV = 4), and $Z_{4}$ (radiotherapy; yes = 1, no = 0). We use the following Cox mediation model to fit this dataset:
@@
Datasets Mediators ${\rm CI_{Sobel}}$ ${\rm CI_{ASobel}}$ $\hat{\alpha}_{k}\hat{\beta}_{k}$
I $M_{1}$ [0.00877, 0.21325] [0.05989, 0.16213] 0.11099
$M_{2}$ [--0.08847, 0.02305] [--0.06058, --0.00483] --0.03271
$M_{3}$ [--0.02641, 0.09471] [0.00387, 0.06443] 0.03416
II $M_{1}$ [--0.00598, 0.02697] [0.00226, 0.01873] 0.01049
III $M_{1}$ [--0.03753, 0.00435] [--0.02707, --0.00612] --0.01659
$M_{2}$ [0.00325, 0.04912] [0.01471, 0.03765] 0.02617
$M_{3}$ [--0.00635, 0.07154] [--0.00635, 0.07154] 0.03260
$M_{4}$ [--0.00363, 0.04097] [0.00752, 0.02982] 0.01868
$M_{5}$ [--0.04995, 0.00811] [--0.04995, 0.00811] --0.02095
$M_{6}$ [--0.06349, 0.00023] [--0.04756, --0.01570] --0.03163
$M_{7}$ [--0.00335, 0.04642] [0.00909, 0.03398] 0.02154
@@ 7 Concluding Remarks
In this paper, we have proposed an data-adjusted joint significance mediation effects test procedure. The explicit expressions of size and power were derived. We also have extended the AJS for performing small-scale multiple testing with FWER control. An adjusted Sobel-type confidence interval was presented. Some simulations and three real-world examples were used to illustrate the usefulness of our method. The proposed method in practical applications is supported by an R package called ”AdapMed”. The focus of this study was limited to a single mediator or a small number of mediators, and the inclusion of high-dimensional mediators was beyond the scope of this research. The following papers are recommended for those interested in high-dimensional mediation analysis, including Zhang et al. (2016; 2021b; 2024), Luo et al. (2020), Perera et al. (2022) and Tian et al. (2022), among others.
There exist two possible directions for applying the proposed AJS test method in our future research. (i) Microbiome Mediation Analysis. Recently, increasing studies have studied the biological mechanisms whether the microbiome play a mediating role between an exposure and a clinical outcome (Sohn and Li, 2019; Wang et al., 2020; Zhang et al., 2021a; Sohn et al., 2022; Yue and Hu, 2022). For improving the powers of mediation effect testing, it is desirable to use the AJS test method when performing microbiome mediation analysis. (ii) Multiple-Mediator Testing with FDR control. We have studied the theoretical and numerical performances of AJS test method for multiple testing with FWER control. Following Sampson et al. (2018), it is useful to investigate the AJS test for multiple mediators with FDR control in some applications.
@@
The Supplementary Material includes Figures S.1-S.3, the proofs of Theorems 1, 2, and 3, the size of ASobel test, together with additional findings for the simulation.
@@ Appendix A Proofs
Proof of Theorem 3. It follows from the definition of ASobel-type confidence interval that
@@
where $T_{max}=\max(|T_{\alpha}|,|T_{\beta}|)$ satisfying $\lim_{n\rightarrow\infty}\mathbb{P}(T_{max}\geq\lambda_{n}|\alpha=\beta=0)=0$ and $\lim_{n\rightarrow\infty}\mathbb{P}(T_{max}<\lambda_{n}|\alpha=\beta=0)=1$. From the definition of ${\rm CI_{ASobel}}$, with probability approaching one, it is straightforward to derive the following expressions:
@@
In a similar way, with probability approaching one, it can be deduced that $\lim_{n\rightarrow\infty}\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta=0)=1-\delta.$
@@
$\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta\neq 0)$
@@
With probability approaching one, it follows from (A), (S.21) and (S.22) that $\lim_{n\rightarrow\infty}\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta\neq 0)=1-\delta.$ This ends the proof.
@@ Appendix B The Size of ASobel Test
In this section, we provide some details on the size of ASobel test. First we deduce the size of our proposed ASobel test method under $H_{00}$. It follows from the decision rule of the ASobel test that
@@
Appendix C Figures S.1-S.3 and Additional Numerical Results
The following section provides Figures S.1-S.3, together with additional findings for the simulation section discussed in the main paper.
@@
Table S.1: The size and power of hypothesis testing with logistic mediation ${}^{\ddagger}$.
$(\alpha,\beta)$ Sobel JS ASobel AJS
$n=500$ (0, 0) 0.0002 0.0023 0.0475 0.0475
(0.5, 0) 0.0455 0.0496 0.0455 0.0496
(0, 0.5) 0.0306 0.0491 0.0582 0.0650
(0.2, 0.2) 0.4227 0.5368 0.5296 0.5900
(0.25, 0.25) 0.6698 0.7241 0.6815 0.7302
$n=1500$ (0, 0) 0.0002 0.0020 0.0508 0.0498
(0.5, 0) 0.0452 0.0460 0.0452 0.0460
(0, 0.5) 0.0439 0.0506 0.0441 0.0507
(0.2, 0.2) 0.9413 0.9493 0.9421 0.9496
(0.25, 0.25) 0.9923 0.9936 0.9923 0.9936
[…]
Table S.2: The size and power of hypothesis testing with Cox mediation ${}^{\ddagger}$.
$(\alpha,\beta)$ Sobel JS ASobel AJS
$n=500$ (0, 0) 0.0004 0.0025 0.0480 0.0472
(0.5, 0) 0.0484 0.0518 0.0484 0.0518
(0, 0.5) 0.0429 0.0487 0.0429 0.0487
(0.15, 0.15) 0.5435 0.7352 0.8281 0.8580
(0.25, 0.25) 0.9934 0.9957 0.9937 0.9957
$n=1500$ (0, 0) 0.0002 0.0022 0.0506 0.0521
(0.5, 0) 0.0537 0.0548 0.0537 0.0548
(0, 0.5) 0.0489 0.0508 0.0489 0.0508
(0.15, 0.15) 0.9955 0.9963 0.9971 0.9977
(0.25, 0.25) 1 1 1 1
[…]
@@
Dimension Sobel JS JT_Comp DACT AJS
$n=500$ $d=10$ FWER 0.0040 0.0084 0.2036 0.0004 0.0400
Power 0.0509 0.1833 0.0362 0.0095 0.3095
$d=15$ FWER 0.0022 0.0050 0.2130 0 0.0390
Power 0.0319 0.1521 0.0392 0.0279 0.2705
$d=20$ FWER 0.0024 0.0054 0.2310 0.0024 0.0536
Power 0.0243 0.1317 0.0489 0.0614 0.2502
$n=1500$ $d=10$ FWER 0.0076 0.0090 0.1990 0 0.0232
Power 0.6040 0.6988 0.2276 0.0022 0.7238
$d=15$ FWER 0.0042 0.0064 0.2152 0 0.0324
Power 0.5477 0.6638 0.2492 0.0184 0.6915
$d=20$ FWER 0.0026 0.0042 0.2382 0 0.0358
Power 0.5075 0.6398 0.2846 0.0907 0.6696
@@
Dimension Sobel JS JT_Comp DACT AJS
$n=500$ $d=10$ FWER 0.0074 0.0126 0.1630 0 0.0286
Power 0.4138 0.5503 0.4234 0.0116 0.6609
$d=15$ FWER 0.0032 0.0056 0.1548 0 0.0320
Power 0.3842 0.5164 0.4219 0.0862 0.6225
$d=20$ FWER 0.0038 0.0054 0.1794 0.0004 0.0384
Power 0.3676 0.4967 0.4325 0.2216 0.6000
$n=1500$ $d=10$ FWER 0.0098 0.0104 0.1618 0 0.0254
Power 0.8475 0.8998 0.6243 0.0006 0.9230
$d=15$ FWER 0.0050 0.0056 0.1678 0 0.0344
Power 0.8131 0.8820 0.6209 0.1289 0.9082
$d=20$ FWER 0.0052 0.0058 0.1954 0 0.0364
Power 0.7922 0.8683 0.6462 0.3465 0.8944
Table S.5: The coverage probability and length of 95% confidence interval with logistic mediation ${}^{\ddagger}$.
@@
$(\alpha_{k},\beta_{k})$ Sobel ASobel Sobel ASobel
$n=500$ (0, 0) 1 0.9532 0.02763 0.01384
(0.35, 0) 0.9544 0.9544 0.17854 0.17854
(0.5, 0) 0.9520 0.9520 0.25399 0.25399
(0, 0.55) 0.9738 0.9038 0.10265 0.09341
(0, 0.75) 0.9620 0.9580 0.13761 0.13700
(0.25, 0.25) 0.9484 0.9398 0.13730 0.13622
(0.35, 0.35) 0.9528 0.9528 0.18783 0.18783
$n=1500$ (0, 0) 1 0.9588 0.00893 0.00447
(0.35, 0) 0.9480 0.9480 0.10026 0.10026
(0.5, 0) 0.9500 0.9500 0.14308 0.14308
(0, 0.55) 0.9612 0.9580 0.05695 0.05672
(0, 0.75) 0.9582 0.9582 0.07716 0.07716
(0.25, 0.25) 0.9484 0.9484 0.07706 0.07706
(0.35, 0.35) 0.9540 0.9540 0.10567 0.10567
@@
Table S.6: The coverage probability and length of 95% confidence interval with Cox mediation ${}^{\ddagger}$.
@@
$(\alpha_{k},\beta_{k})$ Sobel ASobel Sobel ASobel
$n=500$ (0, 0) 1 0.9554 0.01238 0.00620
(0.35, 0) 0.9520 0.9520 0.08082 0.08082
(0.5, 0) 0.9548 0.9548 0.11541 0.11541
(0, 0.5) 0.9588 0.9588 0.08979 0.08979
(0, 0.45) 0.9570 0.9570 0.08050 0.08050
(0.25, 0.25) 0.9404 0.9400 0.07435 0.07420
(0.35, 0.35) 0.9486 0.9486 0.10231 0.10231
$n=1500$ (0, 0) 1 0.9480 0.00413 0.00206
(0.35, 0) 0.9514 0.9514 0.04561 0.04561
(0.5, 0) 0.9448 0.9448 0.06516 0.06516
(0, 0.5) 0.9500 0.9500 0.05104 0.05104
(0, 0.45) 0.9550 0.9550 0.04588 0.04588
(0.25, 0.25) 0.9414 0.9414 0.04184 0.04184
(0.35, 0.35) 0.9510 0.9510 0.05792 0.05792

### v8

@@ 5.3 Coverage Probability of Confidence Interval
The performance of the ASobel-type confidence interval presented in Section 4 is evaluated through simulations conducted in this subsection. The Bootstrap confidence interval and traditional Sobel-type confidence interval, ${\rm CI_{Sobel}}$, given in (4.2), are also considered for comparison. The $100(1-\delta)\%$ Bootstrap confidence interval of mediation effect is given as
${\rm CI_{Bootstrap}}=[Q_{\delta/2}(\{\hat{\alpha}^{(b)}\hat{\beta}^{(b)}\}_{b=1}^{B}),~{}Q_{1-\delta/2}(\{\hat{\alpha}^{(b)}\hat{\beta}^{(b)}\}_{b=1}^{B})],$ (5.1)
where $\hat{\alpha}^{(b)}$ and $\hat{\beta}^{(b)}$ are the corresponding parameter estimators with the $b$-th Bootstrap samples, $b=1,\cdots,B$.
The data are generated in a similar manner as those described in Section 5.2, with the parameters chosen as (i) Linear mediation model : ${\bm{\alpha}}=(0,0.35,0.5,0,0,0.25,0.45)^{\prime}$ and ${\bm{\beta}}=(0,0,0,0.35,0.5,0.35,0.5)^{\prime}$; (ii) Logistic mediation model: ${\bm{\alpha}}=(0,0.35,0.5,0,0,0.35,0.45)^{\prime}$ and ${\bm{\beta}}=(0,0,0,0.75,0.8,0.75,0.85)^{\prime}$; (iii) Cox mediation model: ${\bm{\alpha}}=(0,0.35,0.5,0,0,0.45,0.35)^{\prime}$ and ${\bm{\beta}}=(0,0,0,0.5,0.45,0.55,0.35)^{\prime}$. All the results are based on 5000 repetitions, where $B=1000$, the sample sizes are chosen as $n=200$, 500 and 1000, respectively.
The coverage probability (CP) and length of the 95% confidence interval (LCI) with ${\rm CI_{Sobel}}$, ${\rm CI_{Bootstrap}}$ and ${\rm CI_{ASobel}}$ are reported in Tables 3, S.5 and S.6. The ${\rm CI_{ASobel}}$ outperforms ${\rm CI_{Sobel}}$ in terms of CP under $\alpha=\beta=0$. Additionally, the LCI of ${\rm CI_{ASobel}}$ is significantly shorter compared to that of ${\rm CI_{Sobel}}$ when $\alpha\beta=0$. These findings are consistent with the result of Theorem 3. Lastly, the ASobel confidence interval is much better than Bootstrap method in terms of both CP and LCI in the simulations.
@@
Table 4: The comparison between AJS and AB-JS in terms of size and power.
@@ 6 Real Data Examples
Dataset I: (continuous outcomes). The Louisiana State University Health Sciences Center has explored the relationship between children weight and behavior through a survey of children, teachers and parents in Grenada. The dataset is publicly available within the R package mma. To perform mediation analysis as that of Yu and Li (2017), we set gender as the exposure $X$ (Male =0; Female = 1), and the outcome $Y$ is body mass index (BMI). We consider three mediators: $M_{1}$ (join in a sport team or not), $M_{2}$ (number of hours of exercises per week) and $M_{3}$ (number of hours of sweating activities per week). Furthermore, there are three covariates $Z_{1}$ (age), $Z_{2}$ (number of people in family) and $Z_{3}$ (the number of cars in family). After removing those individuals with missing data, we totally have 646 samples when conducting mediation analysis in the context of linear mediation model. We consider the linear mediation model to fit this dataset:
@@
Dataset III: (time-to-event outcomes). We apply our proposed method to a dataset from The Cancer Genome Atlas (TCGA) lung cancer cohort study, where the data are freely available at https://xenabrowser.net/datapages/. There are 593 patients with non-missing clinical and epigenetic information. From Zhang et al. (2021), we use seven DNA methylation markers as potential mediators: $M_{1}$ (cg02178957), $M_{2}$ (cg08108679), $M_{3}$ (cg21926276), $M_{4}$ (cg26387355), $M_{5}$ (cg24200525), $M_{6}$ (cg07690349) and $M_{7}$ (cg26478297). The exposure $X$ is defined as the number of packs smoked per years, and the survival time is the outcome variable. Two hundred forty three patients died during the follow-up, and the censoring rate is 59%. We are interested in testing the mediation effects of DNA methylation markers along the pathways from smoking to survival of lung cancer patients. Four covariates are included: $Z_{1}$ (age at initial diagnosis), $Z_{2}$ (gender; male = 1, female=0), $Z_{3}$ (tumor stage; Stage I = 1, Stage II = 2, Stage III = 3, Stage IV = 4), and $Z_{4}$ (radiotherapy; yes = 1, no = 0). We use the following Cox mediation model to fit this dataset:
@@
Datasets Mediators ${\rm CI_{Sobel}}$ ${\rm CI_{Bootstrap}}$ ${\rm CI_{ASobel}}$ $\hat{\alpha}_{k}\hat{\beta}_{k}$
I $M_{1}$ [0.0088, 0.2133] [0.0234, 0.2251] [0.0599, 0.1621] 0.1110
$M_{2}$ [-0.0885, 0.0231] [-0.1013, 0.0083] [-0.0606, -0.0048] -0.0327
$M_{3}$ [-0.0264, 0.0947] [-0.0189, 0.1022] [0.0039, 0.0644] 0.0342
II $M_{1}$ [-0.0059, 0.0269] [-0.0022, 0.0335] [0.0023, 0.0187] 0.0105
III $M_{1}$ [-0.0375, 0.0044] [-0.0459, 0.0003] [-0.0271, -0.0061] -0.0166
$M_{2}$ [0.0033, 0.0491] [0.0015, 0.0533] [0.0147, 0.0377] 0.0262
$M_{3}$ [-0.0064, 0.0715] [-0.0062, 0.0835] [-0.0064, 0.0715] 0.0326
$M_{4}$ [-0.0036, 0.0409] [-$10^{-5}$, 0.0471] [0.0075, 0.0298] 0.0187
$M_{5}$ [-0.0499, 0.0081] [-0.0574, 0.0086] [-0.0499, 0.0081] -0.0209
$M_{6}$ [-0.0635, 0.0002] [-0.0733, -0.0044] [-0.0476, -0.0157] -0.0316
$M_{7}$ [-0.0034, 0.0464] [$4\times 10^{-5}$, 0.0519] [0.0091, 0.0339] 0.0215
@@ 7 Concluding Remarks
In this paper, we have proposed an data-adjusted joint significance mediation effects test procedure. The explicit expressions of size and power were derived. We also have extended the AJS for performing small-scale multiple testing with FWER control. An adjusted Sobel-type confidence interval was presented. Some simulations and three real-world examples were used to illustrate the usefulness of our method. The method we propose provides a publicly accessible and user-friendly R package, called $\tt AdjMed$, which can be found at https://github.com/zhxmath/AdjMed. The focus of this study was limited to a single mediator or a small number of mediators, and the inclusion of high-dimensional mediators was beyond the scope of this research.
There exist two possible directions for applying the proposed AJS test method in our future research. (i) Microbiome Mediation Analysis. Recently, increasing studies have studied the biological mechanisms whether the microbiome play a mediating role between an exposure and a clinical outcome (Sohn and Li, 2019). For improving the powers of mediation effect testing, it is desirable to use the AJS test method when performing microbiome mediation analysis. (ii) Multiple-Mediator Testing with FDR control. We have studied the theoretical and numerical performances of AJS test method for multiple testing with FWER control. It is useful to investigate the AJS test for multiple mediators with FDR control in some applications.
@@
The Supplementary Material includes Figures S.1-S.3, Tables S.1-S.6, the proofs of Theorems 1-3, the size of ASobel test, together with the manual for R package AdjMed.
Acknowledgements
We would like to thank the Editor, the Associate Editor and two reviewers for their constructive and insightful comments that greatly improved the manuscript.
@@ Appendix A Proofs
Proof of Theorem 3. By the definition of ASobel-type confidence interval, we have
@@
where the term $T_{max}=\max(|T_{\alpha}|,|T_{\beta}|)$ satisfying $\lim_{n\rightarrow\infty}\mathbb{P}(T_{max}\geq\lambda_{n}|\alpha=\beta=0)=0$ and $\lim_{n\rightarrow\infty}\mathbb{P}(T_{max}<\lambda_{n}|\alpha=\beta=0)=1$. From the definition of ${\rm CI_{ASobel}}$, with probability approaching one, it is straightforward to derive the following expressions:
@@
In a similar way, with probability approaching one, it can be deduced that
$\lim_{n\rightarrow\infty}\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta=0)=1-\delta.$
@@
$\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta\neq 0)$ (S.20)
@@
With probability approaching one, it follows from (S.20), (S.21) and (S.22) that $\lim_{n\rightarrow\infty}\mathbb{P}(\alpha\beta\in{\rm CI_{ASobel}}|\alpha\neq 0,\beta\neq 0)=1-\delta.$ This ends the proof.
@@ Appendix B The Size of ASobel Test
In this section, we provide the proof details for the size of ASobel test. Under $H_{00}$, it follows from the decision rule of the ASobel test that
@@
Appendix C Figures S.1-S.3 and Tables S.1-S.6
The following section provides Figures S.1-S.3, together with Tables S.1-S.6 for the simulation section discussed in the main paper.
@@
Table S.1: The size and power of hypothesis testing with logistic mediation model.
$(\alpha,\beta)$ Sobel JS Bootstrap ASobel AJS
$n=200$ (0, 0) 0 0.0030 0.0024 0.0486 0.0478
(0.5, 0) 0.0368 0.0476 0.0524 0.0368 0.0476
(0, 0.65) 0.0182 0.0486 0.0524 0.0506 0.0724
(0.15, 0.20) 0.0350 0.1354 0.0882 0.2826 0.3206
(0.25, 0.30) 0.2542 0.4296 0.3600 0.3680 0.4974
$n=500$ (0, 0) 0 0.0018 0.0020 0.0476 0.0484
(0.5, 0) 0.0448 0.0472 0.0494 0.0448 0.0472
(0, 0.65) 0.0382 0.0530 0.0570 0.0398 0.0544
(0.15, 0.20) 0.2960 0.4840 0.3982 0.6492 0.6680
(0.25, 0.30) 0.7966 0.8344 0.8354 0.8046 0.8374
$n=1000$ (0, 0) 0 0.0018 0.0010 0.0518 0.0528
(0.5, 0) 0.0516 0.0532 0.0558 0.0516 0.0532
(0, 0.65) 0.0446 0.0510 0.0592 0.0446 0.0510
(0.15, 0.20) 0.7384 0.8124 0.7854 0.8570 0.8708
(0.25, 0.30) 0.9866 0.9886 0.9886 0.9866 0.9886
Table S.2: The size and power of hypothesis testing with Cox mediation model.
$(\alpha,\beta)$ Sobel JS Bootstrap ASobel AJS
$n=200$ (0, 0) 0 0.0030 0.0018 0.0428 0.0448
(0.5, 0) 0.0396 0.0498 0.0494 0.0396 0.0498
(0, 0.5) 0.0376 0.0510 0.0604 0.0382 0.0514
(0.15, 0.15) 0.0802 0.2322 0.2024 0.3552 0.4302
(0.25, 0.25) 0.5738 0.7494 0.7276 0.6522 0.7918
$n=500$ (0, 0) 0.0006 0.0044 0.0028 0.0546 0.0576
(0.5, 0) 0.0512 0.0558 0.0576 0.0512 0.0558
(0, 0.5) 0.0482 0.0570 0.0572 0.0482 0.0570
(0.15, 0.15) 0.5236 0.7266 0.7100 0.8156 0.8452
(0.25, 0.25) 0.9920 0.9946 0.9946 0.9922 0.9948
$n=1000$ (0, 0) 0 0.0026 0.0016 0.0482 0.0480
(0.5, 0) 0.0484 0.0494 0.0520 0.0484 0.0494
(0, 0.5) 0.0432 0.0462 0.0498 0.0432 0.0462
(0.15, 0.15) 0.9486 0.9754 0.9750 0.9786 0.9860
(0.25, 0.25) 1 1 1 1 1
@@
Dimension Sobel JS JT_Comp DACT Bootstrap AJS
$n=200$ $d=10$ FWER 0.0030 0.0058 0.1480 0 0.0092 0.0438
Power 0.0722 0.1586 0.1752 0.0155 0.1645 0.2055
$d=15$ FWER 0.0020 0.0050 0.1568 0.0008 0.0048 0.0474
Power 0.0546 0.1318 0.1721 0.0454 0.1204 0.1676
$d=20$ FWER 0.0016 0.0038 0.1718 0.0020 0.0022 0.0500
Power 0.0401 0.1072 0.1738 0.0740 0.0756 0.1398
$n=500$ $d=10$ FWER 0.0046 0.0094 0.1424 0 0.0136 0.0594
Power 0.3617 0.4860 0.2967 0.0138 0.5004 0.5486
$d=15$ FWER 0.0040 0.0094 0.1576 0 0.0112 0.0592
Power 0.3263 0.4522 0.3012 0.0640 0.4653 0.5116
$d=20$ FWER 0.0026 0.0052 0.1650 0.0004 0.0088 0.0650
Power 0.2998 0.4316 0.3156 0.1651 0.4432 0.4887
$n=1000$ $d=10$ FWER 0.0058 0.0110 0.1432 0 0.0154 0.0468
Power 0.6410 0.7287 0.4536 0.0067 0.7417 0.7677
$d=15$ FWER 0.0040 0.0068 0.1410 0 0.0114 0.0438
Power 0.6018 0.6985 0.4526 0.0771 0.7139 0.7382
$d=20$ FWER 0.0040 0.0050 0.1710 0.0002 0.0092 0.0526
Power 0.5778 0.6809 0.4772 0.2488 0.6993 0.7216
@@
Dimension Sobel JS JT_Comp DACT Bootstrap AJS
$n=200$ $d=10$ FWER 0.0030 0.0108 0.1348 0 0.0124 0.0242
Power 0.0687 0.2009 0.2116 0.0204 0.1713 0.2536
$d=15$ FWER 0.0028 0.0100 0.1480 0.0002 0.0104 0.0334
Power 0.0470 0.1703 0.2049 0.0545 0.1297 0.2128
$d=20$ FWER 0.0030 0.0070 0.1564 0.0036 0.0076 0.0404
Power 0.0349 0.1513 0.2118 0.1037 0.1003 0.1872
$n=500$ $d=10$ FWER 0.0090 0.0120 0.1550 0 0.0148 0.0286
Power 0.4104 0.5449 0.4225 0.0128 0.5412 0.6546
$d=15$ FWER 0.0048 0.0082 0.1566 0 0.0106 0.0370
Power 0.3800 0.5089 0.4175 0.0876 0.5030 0.6148
$d=20$ FWER 0.0038 0.0062 0.1776 0.0004 0.0092 0.0402
Power 0.3614 0.4918 0.4306 0.2160 0.4851 0.5978
$n=1000$ $d=10$ FWER 0.0098 0.0112 0.1666 0 0.0146 0.0254
Power 0.6741 0.7966 0.5256 0.0066 0.8020 0.8470
$d=15$ FWER 0.0060 0.0072 0.1606 0 0.0100 0.0370
Power 0.6330 0.7689 0.5265 0.1056 0.7769 0.8274
$d=20$ FWER 0.0028 0.0034 0.1848 0 0.0078 0.0388
Power 0.6018 0.7490 0.5414 0.3100 0.7589 0.8146
Table S.5: The coverage probability and length of 95% confidence interval with logistic mediation model.
@@
$(\alpha_{k},\beta_{k})$ Sobel Bootstrap ASobel Sobel Bootstrap ASobel
$n=200$ (0, 0) 1 0.9976 0.9462 0.09662 0.18394 0.05037
(0.35, 0) 0.9644 0.9396 0.9596 0.38635 0.60982 0.38482
(0.5, 0) 0.9508 0.9394 0.9508 0.54536 0.85866 0.54536
(0, 0.75) 0.9920 0.9614 0.9030 0.25726 0.38664 0.22188
(0, 0.8) 0.9870 0.9530 0.9072 0.27034 0.40778 0.24024
(0.35, 0.75) 0.9480 0.9182 0.9470 0.48502 0.94437 0.48440
(0.45, 0.85) 0.9526 0.9000 0.9526 0.60453 1.22664 0.60453
$n=500$ (0, 0) 0.9998 0.9976 0.9514 0.03471 0.05171 0.01737
(0.35, 0) 0.9526 0.9416 0.9526 0.22141 0.24178 0.22141
(0.5, 0) 0.9490 0.9450 0.9490 0.31457 0.34101 0.31457
(0, 0.75) 0.9710 0.9482 0.9334 0.14251 0.15533 0.13613
(0, 0.8) 0.9712 0.9502 0.9474 0.15098 0.16404 0.14711
(0.35, 0.75) 0.9492 0.9388 0.9492 0.27651 0.29945 0.27651
(0.45, 0.85) 0.9482 0.9328 0.9482 0.34346 0.37253 0.34346
$n=1000$ (0, 0) 1 0.9994 0.9478 0.01686 0.02423 0.00843
(0.35, 0) 0.9528 0.9454 0.9528 0.15165 0.15741 0.15165
(0.5, 0) 0.9516 0.9496 0.9516 0.21613 0.22400 0.21613
(0, 0.75) 0.9616 0.9472 0.9564 0.09632 0.10023 0.09582
(0, 0.8) 0.9598 0.9454 0.9582 0.10266 0.10656 0.10247
(0.35, 0.75) 0.9422 0.9392 0.9422 0.18969 0.19623 0.18969
(0.45, 0.85) 0.9498 0.9414 0.9498 0.23585 0.24425 0.23585
@@
Table S.6: The coverage probability and length of 95% confidence interval with Cox mediation model.
@@
$(\alpha_{k},\beta_{k})$ Sobel Bootstrap ASobel Sobel Bootstrap ASobel
$n=200$ (0, 0) 0.9998 0.9984 0.9530 0.03376 0.05178 0.01757
(0.35, 0) 0.9624 0.9494 0.9568 0.13767 0.15948 0.13704
(0.5, 0) 0.9488 0.9446 0.9488 0.19484 0.22181 0.19484
(0, 0.5) 0.9688 0.9484 0.9652 0.14885 0.16165 0.14850
(0, 0.45) 0.9746 0.9482 0.9644 0.13393 0.14657 0.13277
(0.45, 0.55) 0.9402 0.9464 0.9402 0.24669 0.27106 0.24669
(0.35, 0.35) 0.9390 0.9500 0.9390 0.17099 0.18929 0.17093
$n=500$ (0, 0) 0.9998 0.9982 0.9490 0.01259 0.01822 0.00632
(0.35, 0) 0.9538 0.9436 0.9538 0.08157 0.08544 0.08157
(0.5, 0) 0.9452 0.9382 0.9452 0.11612 0.12078 0.11612
(0, 0.5) 0.9554 0.9410 0.9554 0.09002 0.09254 0.09002
(0, 0.45) 0.9602 0.9458 0.9602 0.08146 0.08394 0.08146
(0.45, 0.55) 0.9506 0.9458 0.9506 0.14883 0.15322 0.14883
(0.35, 0.35) 0.9490 0.9498 0.9490 0.10299 0.10620 0.10299
$n=1000$ (0, 0) 0.9996 0.9990 0.9504 0.00622 0.00882 0.00311
(0.35, 0) 0.9552 0.9492 0.9552 0.05656 0.05752 0.05656
(0.5, 0) 0.9476 0.9424 0.9476 0.08074 0.08179 0.08074
(0, 0.5) 0.9518 0.9462 0.9518 0.06271 0.06350 0.06271
(0, 0.45) 0.9568 0.9500 0.9568 0.05659 0.05731 0.05659
(0.45, 0.55) 0.9490 0.9448 0.9490 0.10378 0.10486 0.10378
(0.35, 0.35) 0.9506 0.9494 0.9506 0.07145 0.07224 0.07145